OpenStudy (anonymous):

Find the derivative with respect to x and simplify: (3 - 2x)² (x² - 2) Can sum1 help me please? : )

OpenStudy (anonymous):

first simplify them then find the derevative

OpenStudy (anonymous):

it will be easier

OpenStudy (akshay_budhkar):

http://www.wolframalpha.com/input/?i=d%2Fdx%283+-+2x%29%C2%B2+%28x%C2%B2+-+2%29 see here you will get anything just simplify and differentiate

OpenStudy (anonymous):

okay, I would suggest you to treat \[(3-2x)^2 = f(x) \:\:\:\:and\:\:\:\:\:\:x^2-2 = g(x)\]\[Then\:\:\:\:\frac{df(x)g(x)}{dx} = f'(x)g(x) + g'(x)f(x)\]

OpenStudy (anonymous):

16x^36x^2+2x+24

OpenStudy (anonymous):

ishaan is also correct using the product rule

OpenStudy (anonymous):

I see. :)